### v2

# Some explicit values of a $q$-multiple zeta-star function at roots of unity

Abstract.
In this paper, we show some expressions of certain $q$-multiple zeta-star values at roots of unity. These explicit formulas are expressed by using the determinants or Bell polynomials. Explicit formulas for other types of values can be found from recurrence relations obtained using generating functions.
Key words and phrases:
multiple zeta functions, $q$-Stirling numbers with higher level, complete homogeneous symmetric functions, Bell polynomials, determinant
2010 Mathematics Subject Classification:
Primary 11M32; Secondary 05A15, 05A19, 05A30, 11B37, 11B73
1. Introduction
In this paper, we consider the function
(1) $\mathfrak{Z}_{n}^{\star}(q;m,s):=\sum_{1\leq i_{1}\leq i_{2}\leq\dots\leq i_{m}\leq n-1}\frac{1}{(1-q^{i_{1}})^{s}(1-q^{i_{2}})^{s}\cdots(1-q^{i_{m}})^{s}}\,.$
Then we show some fundamental ways to find the values, and give explicit expressions of the values of $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ when $m$ or $s$ are comparatively small. When $n\to\infty$, the values in (1) may be called one of the $q$-multiple zeta-star values (see, e.g., [17, 19]).
In previous papers ([8, 10]), as an application of the study of certain generalized Stirling numbers ([7, 8]), we considered a function
(2) $\mathfrak{Z}_{n}(q;m,s):=\sum_{1\leq i_{1}<i_{2}<\dots<i_{m}\leq n-1}\frac{1}{(1-q^{i_{1}})^{s}(1-q^{i_{2}})^{s}\cdots(1-q^{i_{m}})^{s}}\,.$
When $n\to\infty$, these values may be called one of the $q$-multiple zeta values (see, e.g., [3, 18, 20, 22, 23]). Then when $q$ takes the root of unity $\zeta_{n}$ in (2), some explicit expressions can be given:
(3) $\mathfrak{Z}_{n}(\zeta_{n};m,1)$ $=\frac{1}{m+1}\binom{n-1}{m}\,,$
(4) $\mathfrak{Z}_{n}(\zeta_{n};m,2)$ $=\frac{1}{n(m+1)}\left(\binom{n-1}{m}+(-1)^{m}\binom{n-1}{2m+1}\right)$
(5) $=\frac{2\cdot m!}{(2m+2)!}\binom{n-1}{m}\sum_{k=0}^{m}\left[\!\!\left[2m+2\atop m+k+2\right]\!\!\right]_{1}^{(m+1,1)}(-n)^{k}\,,$
where $\left[\!\!\left[n\atop k\right]\!\!\right]_{1}^{(r,1)}$ denotes the $r$-Stirling number ([4]). We also have
(6) $\mathfrak{Z}_{n}(\zeta_{n};1,s)=\left|\begin{array}[]{ccccc}\frac{n-1}{2}&1&0&\cdots&\\ \frac{2}{3}\binom{n-1}{2}&\frac{n-1}{2}&1&&\vdots\\ \vdots&&\ddots&&0\\ \frac{s-1}{s}\binom{n-1}{s-1}&\frac{1}{s-1}\binom{n-1}{s-2}&\cdots&\frac{n-1}{2}&1\\ \frac{s}{s+1}\binom{n-1}{s}&\frac{1}{s}\binom{n-1}{s-1}&\cdots&\frac{1}{3}\binom{n-1}{2}&\frac{n-1}{2}\\ \end{array}\right|\,.$
An expression of $\mathfrak{Z}_{n}(\zeta_{n};m,s)$ ($s=3,4,\dots$) would be possible, but it looks to become very complicated.
In addition, even though there is no direct expression of $\mathfrak{Z}_{n}(\zeta_{n};m,s)$ ($m=2,3,\dots$), its value may be calculated by using the known ones.
For example, when $m=2$, we can obtain $\mathfrak{Z}_{n}(\zeta_{n};2,s)$ as
(7) $\mathfrak{Z}_{n}(\zeta_{n};2,s)=\frac{1}{2}\left(\mathfrak{Z}_{n}(\zeta_{n};1,s)^{2}-\mathfrak{Z}_{n}(\zeta_{n};1,2s)\right)\,.$
Similarly, for $m=3$,
(8) $\mathfrak{Z}_{n}(\zeta_{n};3,s)=\frac{1}{6}\left(\mathfrak{Z}_{n}(\zeta_{n};1,s)^{3}-3\mathfrak{Z}_{n}(\zeta_{n};1,s)\mathfrak{Z}_{n}(\zeta_{n};1,2s)+2\mathfrak{Z}_{n}(\zeta_{n};1,3s)\right)\,.$
It is noticed that, while many explicit formulas for the $q$-multiple zeta values are known, few explicit formulas for the $q$-multiple zeta-star values are known, even though the relationships involving the $q$-multiple zeta-star values are known (see, e.g., [1, 2, 14]). One of the reasons is that it is more difficult to handle the complete homogeneous symmetric functions corresponding to the $q$-multiple zeta-star values than the elementary symmetric functions corresponding to the $q$-multiple zeta values.
2. $q$-generalized $(r,s)$-Stirling numbers
In [8], $q$-generalized $(r,s)$-Stirling numbers of both kinds are introduced. These Stirling numbers are closely related to our $q$-multiple zeta function in (2) and $q$-multiple zeta-star function in (1).
Let $[n]_{q}$ denote the $q$-number, defined by
$[n]_{q}=\frac{q^{n}-1}{q-1}\quad(q\neq 1)\,.$
Its $q$-factorial is given by $[n]_{q}!=[n]_{q}[n-1]_{q}\cdots[1]_{q}$ with $[0]_{q}!=1$. Let $r$ be a positive integer. The $q$-version of $r$-Stirling numbers of the first kind with higher level (level $s$) are denoted by $\left[\!\!\left[n\atop k\right]\!\!\right]_{q}^{(r,s)}$, and appear in the coefficients in the expansion
(9) $(x)_{n,q}^{(r,s)}=\sum_{k=0}^{n}(-1)^{n-k}\left[\!\!\left[n\atop k\right]\!\!\right]_{q}^{(r,s)}x^{k}\,,$
where for $r,s\geq 1$, $(x)_{n,q}^{(r,s)}$ is defined by
$(x)_{n,q}^{(r,s)}:=x^{r}\prod_{i=r}^{n-1}\bigl{(}x-([i]_{q})^{s}\bigr{)}\quad(n>r)$
with $(x)_{r,q}^{(r,s)}=x^{r}$. When $r=s=1$, $s_{q}(n,k)=(-1)^{n-k}\left[\!\!\left[n\atop k\right]\!\!\right]_{q}^{(1,1)}$ are the signed $q$-Stirling numbers of the first kind (see, e.g., [5]), and $\left[\!\!\left[n\atop k\right]\!\!\right]_{q}=\left[\!\!\left[n\atop k\right]\!\!\right]_{q}^{(1,1)}$ are the unsigned $q$-Stirling numbers of the first kind. When $r=1$ and $q\to 1$, $\left[\!\!\left[n\atop k\right]\!\!\right]^{(s)}=\left[\!\!\left[n\atop k\right]\!\!\right]_{1}^{(1,s)}$ are the (unsigned) Stirling numbers of the first kind with higher level ([12]). When $r=s=1$ and $q\to 1$, $\left[n\atop k\right]=\left[\!\!\left[n\atop k\right]\!\!\right]_{1}^{(1,1)}$ are the unsigned Stirling numbers of the first kind.
The $q$-version of $r$-Stirling numbers of the second kind with higher level are denoted by $\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(r,s)}$, and appear in the coefficients in the expansion
(10) $x^{n}=\sum_{k=0}^{n}\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(r,s)}(x)_{k,q}^{(r,s)}\,.$
When $r=1$ and $q\to 1$, $\left\{\!\!\left\{n\atop k\right\}\!\!\right\}^{(s)}=\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{1}^{(1,s)}$ are the Stirling numbers of the second kind with higher level, studied in [13]. When $r=s=1$, $S_{q}(n,k)=\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}=(-1)^{n-k}\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(1,1)}$ are the signed $q$-Stirling numbers of the second kind (see, e.g., [5]). When $r=s=1$ and $q\to 1$, $\left\{n\atop k\right\}=\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{1}^{(1,1)}$ are the classical Stirling numbers of the second kind.
From the definitions in (9) and (10), the following recurrence relations hold.
(11) $\left[\!\!\left[n\atop k\right]\!\!\right]_{q}^{(r,s)}=\left[\!\!\left[n-1\atop k-1\right]\!\!\right]_{q}^{(r,s)}+\bigl{(}[n-1]_{q}\bigr{)}^{s}\left[\!\!\left[n-1\atop k\right]\!\!\right]_{q}^{(r,s)}$
with
$\left[\!\!\left[n\atop k\right]\!\!\right]_{q}^{(r,s)}=0\quad(0\leq k\leq r,\,n\geq k),$
$\left[\!\!\left[n\atop k\right]\!\!\right]_{q}^{(r,s)}=0\quad(n<k),\quad\left[\!\!\left[0\atop 0\right]\!\!\right]_{q}^{(r,s)}=1\,.$
(12) $\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(r,s)}=\left\{\!\!\left\{n-1\atop k-1\right\}\!\!\right\}_{q}^{(r,s)}+\bigl{(}[k]_{q}\bigr{)}^{s}\left\{\!\!\left\{n-1\atop k\right\}\!\!\right\}_{q}^{(r,s)}$
with
$\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(r,s)}=0\quad(0\leq k\leq r-1,~{}n\geq k),\quad\left\{\!\!\left\{0\atop 0\right\}\!\!\right\}_{q}^{(r,s)}=1,$
$\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(r,s)}=0\quad(n\leq k)\,.$
By the definitions and/or the recurrence relations, we can get expressions with combinatorial summations ([8]).
Lemma 1.
For $r\leq m\leq n-1$ and $r\geq 1$, we have
$\left[\!\!\left[n\atop m\right]\!\!\right]_{q}^{(r,s)}=\left(\frac{[n-1]_{q}!}{[r-1]_{q}!}\right)^{s}\sum_{r\leq i_{1}<\dots<i_{m-r}\leq n-1}\frac{1}{\bigl{(}[i_{1}]_{q}\dots[i_{m-r}]_{q}\bigr{)}^{s}}\,.$
For $n-m\geq r$ and $r\geq 1$, we have
$\left[\!\!\left[n\atop n-m\right]\!\!\right]_{q}^{(r,s)}$ $=\sum_{r\leq i_{1}<\dots<i_{m}\leq n-1}\bigl{(}[i_{1}]_{q}\dots[i_{m}]_{q}\bigr{)}^{s}$
$=\sum_{r\leq i_{1}\leq i_{2}\leq\dots\leq i_{m}\leq n-m}([i_{1}]_{q}[i_{2}+1]_{q}\cdots[i_{m}+m-1]_{q})^{s}$
$=\sum_{i_{m}=r}^{n-m}([i_{m}+m-1]_{q})^{s}\sum_{i_{m-1}=r}^{i_{m}}([i_{m-1}+m-2]_{q})^{s}\cdots\sum_{i_{2}=r}^{i_{3}}([i_{2}+1]_{q})^{s}\sum_{i_{1}=r}^{i_{2}}([i_{1}]_{q})^{s}\,.$
Lemma 2.
For $r+1\leq k\leq n$ and $r\geq 1$,
$\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(r,s)}=\sum_{i_{k-r}=0}^{n-k}\bigl{(}[k]_{q}\bigr{)}^{(n-k-i_{k-r})s}\sum_{i_{k-r-1}=0}^{i_{k-r}}\bigl{(}[k-1]_{q}\bigr{)}^{(i_{k-r}-i_{k-r-1})s}\\ \cdots\sum_{i_{2}=0}^{i_{3}}\bigl{(}[r+2]_{q}\bigr{)}^{(i_{3}-i_{2})s}\sum_{i_{1}=0}^{i_{2}}\bigl{(}[r+1]_{q}\bigr{)}^{(i_{2}-i_{1})s}\bigl{(}[r]_{q}\bigr{)}^{i_{1}s}\,.$
For $n-k\geq r\geq 1$,
$\left\{\!\!\left\{n\atop n-k\right\}\!\!\right\}_{q}^{(r,s)}=\sum_{r\leq i_{1}\leq i_{2}\leq\cdots\leq i_{k}\leq n-k}([i_{1}]_{q}[i_{2}]_{q}\cdots[i_{k}]_{q})^{s}\,.$
From Lemma 1 and Lemma 2, $q$-multiple zeta functions and $q$-multiple zeta-star functions can be written in terms of $q$-generalized $(1,s)$-Stirling numbers of the first kind and of the second kind, respectively.
Proposition 1.
For positive integers $n$, $m$ and $s$, we have
$\mathfrak{Z}_{n}(q;m,s)$ $=\frac{1}{(1-q)^{ms}\bigl{(}[n-1]_{q}!\bigr{)}^{s}}\left[\!\!\left[n\atop m+1\right]\!\!\right]_{q}^{(1,s)}\,,$
$\mathfrak{Z}_{n}^{\star}(q;m,s)$ $=\frac{1}{(1-q)^{ms}}\left\{\!\!\left\{n+m-1\atop n-1\right\}\!\!\right\}_{q}^{(1,-s)}\,.$
3. The main result
Note that $\mathfrak{Z}_{n}^{\star}(q;1,s)=\mathfrak{Z}_{n}(q;1,s)$. Hence, one can have interest in the case where $m\geq 2$. Then, in light of the result of $\mathfrak{Z}_{n}(q;m,s)$, when $q$ takes the root of unity, we can get the following determinant expression.
Theorem 1.
For integers $n,m$ with $n\geq 2$ and $m\geq 1$, we have
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$
$=\frac{1}{m!}\mathbf{Y}_{m}\bigl{(}\mathfrak{Z}_{n}(\zeta_{n};1,s),1!\mathfrak{Z}_{n}(\zeta_{n};1,2s),2!\mathfrak{Z}_{n}(\zeta_{n};1,3s),3!\mathfrak{Z}_{n}(\zeta_{n};1,4s),\dots)\bigr{)}$
$=\frac{1}{m!}\left|\begin{array}[]{ccccc}\mathfrak{Z}_{n}(\zeta_{n};1,s)&-1&0&\cdots&\\ \mathfrak{Z}_{n}(\zeta_{n};1,2s)&\mathfrak{Z}_{n}(\zeta_{n};1,s)&-2&&\vdots\\ \vdots&&\ddots&&0\\ \mathfrak{Z}_{n}\bigl{(}\zeta_{n};1,(m-1)s\bigr{)}&\mathfrak{Z}_{n}\bigl{(}\zeta_{n};1,(m-2)s\bigr{)}&\cdots&\mathfrak{Z}_{n}(\zeta_{n};1,s)&-m+1\\ \mathfrak{Z}_{n}(\zeta_{n};1,ms)&\mathfrak{Z}_{n}\bigl{(}\zeta_{n};1,(m-1)s\bigr{)}&\cdots&\mathfrak{Z}_{n}(\zeta_{n};1,2s)&\mathfrak{Z}_{n}(\zeta_{n};1,s)\\ \end{array}\right|\,.$
For example, when $m=2$, we have
(13) $\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,s)=\frac{1}{2}\left(\mathfrak{Z}_{n}(\zeta_{n};1,s)^{2}+\mathfrak{Z}_{n}(\zeta_{n};1,2s)\right)\,.$
Another way is
(14) $\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,s)=\mathfrak{Z}_{n}(\zeta_{n};2,s)+\mathfrak{Z}_{n}(\zeta_{n};1,2s)\,.$
Then for $s=1,2,3$ we have
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,1)$ $=\frac{(n+1)(n-1)}{12}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,2)$ $=\frac{(n+1)(n-1)(n-3)(n-7)}{2\cdot 5!}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,3)$ $=-\frac{(n+1)(n-1)(n^{4}-650n^{2}+3780n-5291)}{12\cdot 7!}\,.$
For example, when $m=2$ and $s=3$, by using the known values of $\mathfrak{Z}_{n}(\zeta_{n};2,3)$ (or $\mathfrak{Z}_{n}(\zeta_{n};1,3)$) and $\mathfrak{Z}_{n}(\zeta_{n};1,6)$, we have
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,3)=\frac{1}{2}\left(\mathfrak{Z}_{n}(\zeta_{n};1,3)^{2}+\mathfrak{Z}_{n}(\zeta_{n};1,6)\right)$
$=\frac{1}{2}\left(\frac{(n-1)^{2}(n-3)^{2}}{8^{2}}-\frac{(n-1)(2n^{5}+2n^{4}-355n^{3}-355n^{2}+11153n-19087)}{12\cdot 7!}\right)$
$=-\frac{(n+1)(n-1)(n^{4}-650n^{2}+3780n-5291)}{12\cdot 7!}$
or
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,3)=\mathfrak{Z}_{n}(\zeta_{n};2,3)+\mathfrak{Z}_{n}(\zeta_{n};1,6)$
$=\frac{6(n-1)(n-2)(n^{4}+3n^{3}+301n^{2}-2883n+6898)}{9!}$
$\quad-\frac{(n-1)(2n^{5}+2n^{4}-355n^{3}-355n^{2}+11153n-19087)}{12\cdot 7!}$
$=-\frac{(n+1)(n-1)(n^{4}-650n^{2}+3780n-5291)}{12\cdot 7!}\,.$
In order to prove Theorem 1, we need the following Lemmas ([6, Proposition 1],[15, Ch.I§2]). Notice that in the case of (2), the equivalent of elementary symmetric functions arises, whereas in the case of (1), the equivalent of complete homogeneous symmetric functions arises.
Lemma 3.
Let $n$, $M$ and $K$ be positive integers with $M\geq K$. For $g_{n}:=a_{1}^{n}+a_{2}^{n}+\cdots+a_{M}^{n}$ we have
$\sum_{1\leq j_{1}\leq j_{2}\leq\dots\leq j_{K}\leq M}a_{j_{1}}a_{j_{2}}\cdots a_{j_{K}}=\frac{1}{K!}\mathbf{Y}_{K}(g_{1},1!g_{2},2!g_{3},3!g_{4},\dots)\,,$
where $\mathbf{Y}_{n}(x_{1},x_{2},\dots,x_{n})$ is the (complete exponential) Bell polynomial, defined by
$\exp\left(\sum_{m=1}^{\infty}x_{m}\frac{t^{m}}{m!}\right)=1+\sum_{n=1}^{\infty}\mathbf{Y}_{n}(x_{1},x_{2},\dots,x_{n})\frac{t^{n}}{n!}\,.$
That is,
$\mathbf{Y}_{n}(x_{1},x_{2},x_{3},\dots,x_{n})$
$=\sum_{k=1}^{n}\sum_{i_{1}+2i_{2}+\cdots+(n-k+1)i_{n-k+1}=n\atop i_{1}+i_{2}+i_{3}+\cdots=k}\frac{n!}{i_{1}!i_{2}!\cdots i_{n-k+1}!}\left(\frac{x_{1}}{1!}\right)^{i_{1}}\left(\frac{x_{2}}{2!}\right)^{i_{2}}\cdots\left(\frac{x_{n-k+1}}{(n-k+1)!}\right)^{i_{n-k+1}}$
with $\mathbf{Y}_{0}=1$.
Lemma 4.
The following expressions are equivalent.
  1. (i)
    $b_{m}=\sum_{i_{1}+2i_{2}+\cdots+mi_{m}=m\atop i_{1},i_{2},\dots,i_{m}\geq 0}\frac{1}{i_{1}!i_{2}!\cdots i_{m}!}\left(\frac{a_{1}}{1}\right)^{i_{1}}\left(\frac{a_{2}}{2}\right)^{i_{2}}\cdots\left(\frac{a_{m}}{m}\right)^{i_{m}}$
  2. (ii)
    $b_{m}=\frac{1}{m!}\left|\begin{array}[]{ccccc}a_{1}&-1&0&\cdots&0\\ a_{2}&a_{1}&-2&&\vdots\\ \vdots&&\ddots&&0\\ a_{m-1}&a_{m-2}&\cdots&a_{1}&-m+1\\ a_{m}&a_{m-1}&\cdots&a_{2}&a_{1}\\ \end{array}\right|$
  3. (iii
    $a_{n}=(-1)^{n-1}\left|\begin{array}[]{ccccc}b_{1}&1&0&\cdots&0\\ 2b_{2}&b_{1}&1&&\vdots\\ \vdots&&\ddots&&0\\ (n-1)b_{n-1}&b_{n-2}&\cdots&b_{1}&1\\ nb_{n}&b_{n-1}&\cdots&b_{2}&b_{1}\\ \end{array}\right|$
  4. (iv)
    $mb_{m}=\sum_{i=1}^{m}a_{i}b_{m-i}$
  5. (v)
    $a_{n}=-\sum_{j=1}^{n-1}b_{j}a_{n-j}+nb_{n}$
Remark. The relations between (i) and (ii) are yielded from a simple modification of Trudi’s formula ([16, Vol.3, p.214],[21]). By applying the inversion formula (see, e.g., [9, Lemma 1],[11, Theorem 1]), we can find that (iv) and (v) are equivalent. Others can be yielded by the expansions of the determinants.
Proof of Theorem 1..
Put $b_{m}=\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ and $a_{j}=\mathfrak{Z}_{n}(\zeta_{n};1,js)$ in Lemma 4 by applying Lemma 3. ∎
4. The values $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ for small $s$
In the previous section, we considered the values $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ for small $m$. In this section we shall consider the values $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ for small $s$.
$\mathfrak{Z}_{n}(\zeta_{n};m,1)$ is expressed in a simple form as in (3), but $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,1)$ is not.
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};1,1)$ $=\frac{n-1}{2}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,1)$ $=\frac{(n+1)(n-1)}{12}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};3,1)$ $=\frac{(n+1)(n-1)}{24}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};4,1)$ $=-\frac{(n+1)(n-1)(n^{2}-19)}{720}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};5,1)$ $=-\frac{(n+1)(n-1)(n+3)(n-3)}{480}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};6,1)$ $=\frac{(n+1)(n-1)(2n^{4}-145n^{2}+863)}{60480}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};7,1)$ $=\frac{(n+1)(n-1)(n+5)(n-5)(2n^{2}-11)}{24192}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};8,1)$ $=-\frac{(n+1)(n-1)(3n^{6}-497n^{4}+9247n^{2}-33953)}{3628800}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};9,1)$ $=-\frac{(n+1)(n-1)(n+7)(n-7)(3n^{4}-50n^{2}+167)}{1036800}\,.$
As $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ is a complete homogeneous symmetric polynomial, let us consider the generating function:
$\sum_{m=0}^{n-1}\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)X^{m}=\sum_{m=0}^{\infty}\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)X^{m}$
$=\sum_{m=0}^{\infty}\sum_{1\leq i_{1}\leq i_{2}\leq\dots\leq i_{m}\leq n-1}\frac{X^{m}}{(1-\zeta_{n}^{i_{1}})^{s}(1-\zeta_{n}^{i_{2}})^{s}\cdots(1-\zeta_{n}^{i_{m}})^{s}}$
$=\prod_{j=1}^{n-1}\dfrac{1}{1-\dfrac{X}{(1-\zeta_{n}^{j})^{s}}}=n^{s}\prod_{j=1}^{n-1}\frac{1}{(1-\zeta_{n}^{j})^{s}-X}\,.$
Let $\alpha_{i}$ ($i=1,2,\dots,s$) be the roots of the polynomial $(1-Y)^{s}-X$. Namely,
(15) $\prod_{i=1}^{s}(\alpha_{i}-Y)=(1-Y)^{s}-X\,.$
Since by (15)
$\prod_{i=1}^{s}(\alpha_{i}-1)=-X\,,$
we get
$\prod_{j=1}^{n-1}\prod_{i=1}^{s}(\alpha_{i}-\zeta_{n}^{j})=\prod_{i=1}^{s}\frac{\alpha_{i}^{n}-1}{\alpha_{i}-1}=\frac{-1}{X}\prod_{i=1}^{s}(\alpha_{i}^{n}-1)\,.$
Hence,
(16) $\sum_{m=0}^{n-1}\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)X^{m}=-Xn^{s}\prod_{i=1}^{s}\frac{1}{\alpha_{i}^{n}-1}\,.$
4.1. The values $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,1)$
Let us consider the value $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ when $s=1$. When $s=1,2$, it is simpler and faster to calculate the values directly, though there is a general method that is described in Subsection 4.3.
When $s=1$, by $\alpha_{1}-Y=(1-Y)-X$ in (15), we see that $\alpha_{1}=1-X$. Thus, by (16) we have
$\sum_{m=0}^{n-1}\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,1)X^{m}$ $=n\prod_{j=1}^{n-1}\frac{1}{(1-\zeta_{n}^{j})-X}$
$=\frac{-Xn}{\alpha_{1}^{n}-1}=\frac{Xn}{1-(1-X)^{n}}$
$=\frac{n}{\sum_{\kappa=0}^{n-1}(1-X)^{\kappa}}=\frac{n}{\sum_{k=0}^{n-1}\binom{n}{k+1}(-1)^{k}X^{k}}$
$=\frac{1}{\sum_{k=0}^{n-1}\binom{n-1}{k}\frac{(-1)^{k}}{k+1}X^{k}}=\frac{1}{1-\sum_{k=1}^{n-1}\binom{n-1}{k}\frac{(-1)^{k+1}}{k+1}X^{k}}$
Hence,
11 11 $=\left(\sum_{k=0}^{n-1}\binom{n-1}{k}\frac{(-1)^{k}}{k+1}X^{k}\right)\left(\sum_{m=0}^{n-1}\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,1)X^{m}\right)$
$=\sum_{l=0}^{n-1}\left(\sum_{j=0}^{l}\frac{(-1)^{l-j}}{l-j+1}\binom{n-1}{l-j}\mathfrak{Z}_{n}^{\star}(\zeta_{n};j,1)\right)X^{l}$
$\quad+\sum_{l=n}^{2n-2}\left(\sum_{j=l-n+1}^{n-1}\frac{(-1)^{l-j}}{l-j+1}\binom{n-1}{l-j}\mathfrak{Z}_{n}^{\star}(\zeta_{n};j,1)\right)X^{l}\,.$
Comparing the coefficients on both sides, we have the recurrence formula when $s=1$. By applying this formula repeatedly, we can get explicit expressions of $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,1)$ ($m=1,2,\dots,9$) as seen above.
Theorem 2.
For $11\leq m\leq n-$ 1, we have
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,1)=\sum_{j=0}^{m-1}\frac{(-1)^{m-j+1}}{m-j+1}\binom{n-1}{m-j}\mathfrak{Z}_{n}^{\star}(\zeta_{n};j,1)$
with $\mathfrak{Z}_{n}^{\star}(\zeta_{n};0,1)=1$.
4.2. The values $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,2)$
Let us consider the value $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ when $s=2$. When $s=2$, by (15), we see that $\alpha_{1}+\alpha_{2}=2$ and $\alpha_{1}\alpha_{2}=1-X$. Thus, by (16) we have
$\sum_{m=0}^{n-1}\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,2)X^{m}$ $=n^{2}\prod_{j=1}^{n-1}\frac{1}{(1-\zeta_{n}^{j})^{2}-X}$
$=\frac{-Xn^{2}}{(\alpha_{1}^{n}-1)(\alpha_{2}^{n}-1)}=\frac{-Xn^{2}}{1-\bigl{(}(1+\sqrt{X})^{n}+(1-\sqrt{X})^{n}\bigr{)}+(1-X)^{n}}$
$=\frac{-Xn^{2}}{1-2\sum_{\kappa=0}^{\left\lfloor\frac{n}{2}\right\rfloor}\binom{n}{2\kappa}X^{\kappa}+\sum_{\kappa=0}^{n}\binom{n}{\kappa}(-X)^{\kappa}}$
$=\frac{-n^{2}}{-2\sum_{\kappa=1}^{\left\lfloor\frac{n}{2}\right\rfloor}\binom{n}{2\kappa}X^{\kappa-1}+\sum_{\kappa=1}^{n}\binom{n}{\kappa}(-1)^{\kappa}X^{\kappa-1}}$
$=\frac{n^{2}}{2\sum_{\kappa=0}^{\left\lfloor\frac{n}{2}\right\rfloor-1}\binom{n}{2\kappa+2}X^{\kappa}+\sum_{\kappa=0}^{n-1}\binom{n}{\kappa+1}(-1)^{\kappa}X^{\kappa}}\,.$
Hence,
11 11 $=\left(\frac{2}{n^{2}}\sum_{\kappa=0}^{\left\lfloor\frac{n}{2}\right\rfloor-1}\binom{n}{2\kappa+2}X^{\kappa}\right)\left(\sum_{m=0}^{n-1}\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,2)X^{m}\right)$
$\quad+\left(\frac{1}{n^{2}}\sum_{\kappa=0}^{n-1}\binom{n}{\kappa+1}(-1)^{\kappa}X^{\kappa}\right)\left(\sum_{m=0}^{n-1}\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,2)X^{m}\right)$
$=\sum_{l=0}^{\left\lfloor\frac{n}{2}\right\rfloor-1}\left(\sum_{j=0}^{l}\frac{2}{n^{2}}\binom{n}{2l-2j+2}\mathfrak{Z}_{n}^{\star}(\zeta_{n};j,2)\right)X^{l}$
$\quad+\sum_{l=0}^{n-1}\left(\sum_{j=0}^{l}\frac{(-1)^{l-j}}{n^{2}}\binom{n}{l-j+1}\mathfrak{Z}_{n}^{\star}(\zeta_{n};j,2)\right)X^{l}$
$\quad+R(n,l)\,,$
where $R(n,l)$ implies the terms for $l\geq\left\lfloor\frac{n}{2}\right\rfloor$ and for $l\geq n$, respectively. Comparing the coefficients on both sides, we have for $11\leq l\leq\left\lfloor\frac{n}{2}\right\rfloor-$ 1
$\sum_{j=0}^{l}\frac{1}{n^{2}}\left(2\binom{n}{2l-2j+2}+(-1)^{l-j}\binom{n}{l-j+1}\right)\mathfrak{Z}_{n}^{\star}(\zeta_{n};j,2)=0\,.$
Hence, we have the following.
Theorem 3.
For $11\leq m\leq\left\lfloor\frac{n}{2}\right\rfloor-$ 1, we have
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,2)=-\sum_{j=0}^{m-1}\frac{1}{n^{2}}\left(2\binom{n}{2m-2j+2}+(-1)^{m-j}\binom{n}{m-j+1}\right)\mathfrak{Z}_{n}^{\star}(\zeta_{n};j,2)$
with $\mathfrak{Z}_{n}^{\star}(\zeta_{n};0,2)=1$.
By applying this formula repeatedly, we can get explicit expressions of $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,2)$ ($m=1,2,3,4,5$) as follows.
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};1,2)$ $=-\frac{(n-1)(n-5)}{12}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,2)$ $=\frac{(n+1)(n-1)(n-3)(n-7)}{240}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};3,2)$ $=-\frac{(n+1)(n-1)(10n^{4}-126n^{3}+241n^{2}+1134n-2699)}{12\cdot 7!}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};4,2)$ $=\frac{(n+1)(n-1)(21n^{6}-300n^{5}+181n^{4}+9150n^{3}-21071n^{2}-41250n+103669)}{10!}$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};5,2)$ $=-\frac{(n+1)(n-1)(n-3)}{4\cdot 11!}\times(30n^{7}-372n^{6}-1889n^{5}$
$\qquad+24671n^{4}+18790n^{3}-346648n^{2}-57251n+1088429)\,.$
4.3. The cases for $s\geq 3$
It is not easy to find an explicit expression of $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ when $s\geq 3$ though it is possible theoretically.
The following method was hinted by Henrik Bachmann directly when the author visited his university in 2024 ([1, 2]). In order to calculate $\prod_{i=1}^{s}(\alpha_{i}^{n}-1)^{-1}$, we consider the function
$f(X,Y):=\sum_{n=1}^{\infty}\frac{Y^{n}}{n}\prod_{i=1}^{s}(\alpha_{i}^{n}-1)$
$=\sum_{n=1}^{\infty}\frac{Y^{n}}{n}\sum_{l=0}^{s}(-1)^{s-l}\sum_{1\leq i_{1}<\cdots<i_{l}\leq s}(\alpha_{i_{1}}\cdots\alpha_{i_{l}})^{n}$
$=(-1)^{s-1}\sum_{l=0}^{s}(-1)^{l-1}\sum_{1\leq i_{1}<\cdots<i_{l}\leq s}\sum_{n=1}^{\infty}\frac{(\alpha_{i_{1}}\cdots\alpha_{i_{l}}Y)^{n}}{n}$
$=(-1)^{s-1}\sum_{l=0}^{s}(-1)^{l}\sum_{1\leq i_{1}<\cdots<i_{l}\leq s}\log(1-\alpha_{i_{1}}\cdots\alpha_{i_{l}}Y)$
$=(-1)^{s-1}\log\left(\prod_{l=0}^{s}F_{s,l}(X,Y)^{(-1)^{l}}\right)\,,$
where
(17) $F_{s,l}(X,Y):=\prod_{1\leq i_{1}<\cdots<i_{l}\leq s}(1-\alpha_{i_{1}}\cdots\alpha_{i_{l}}Y)\quad(1\leq l\leq s)$
with $F_{s,0}(X,Y)=1-Y$. Here, $F_{s,l}(X,Y)$’s include the elementary symmetric functions but can be determined from (15) as the complete homogeneous symmetric functions.
For example, we can find that
$F_{1,1}(X,Y)$ $=1-Y+XY\,,$
$F_{2,1}(X,Y)$ $=(1-Y)^{2}-XY^{2}\,,$
$F_{2,2}(X,Y)$ $=1-Y+XY\,,$
$F_{3,1}(X,Y)$ $=(1-Y)^{3}+XY^{3}\,,$
$F_{3,2}(X,Y)$ $=(1-Y)^{3}-3XY^{2}-(X^{2}-2X)Y^{3}\,,$
$F_{3,3}(X,Y)$ $=1-Y+XY\,.$
Though it becomes more complicated, using $F_{1,1}(X,Y)$, or $F_{2,1}(X,Y)$ and $F_{2,2}(X,Y)$, we can find the expression $\prod_{i=1}^{s}(\alpha_{i}^{n}-1)$ for $s=1$ or $s=2$, respectively.
Let us find the expression $\prod_{i=1}^{s}(\alpha_{i}^{n}-1)$ for $s=3$ by using $F_{3,1}(X,Y)$, $F_{3,2}(X,Y)$ and $F_{3,3}(X,Y)$.
$f(X,Y)=(-1)^{3-1}\log\left(\prod_{l=0}^{3}F_{3,l}(X,Y)^{(-1)^{l}}\right)$
$=\log\frac{(1-Y)\bigl{(}(1-Y)^{3}-3XY^{2}-(X^{2}-2X)Y^{3}\bigr{)}}{\bigl{(}(1-Y)^{3}+XY^{3}\bigr{)}(1-Y+XY)}$
$=\log\left(1-\frac{3XY^{2}+(X^{2}-2X)Y^{3}}{(1-Y)^{3}}\right)-\log\left(1+\frac{XY^{3}}{(1-Y)^{3}}\right)-\log\left(1+\frac{XY}{1-Y}\right)$
$=-\sum_{n=1}^{\infty}\frac{1}{n}\left(\frac{3XY^{2}+(X^{2}-2X)Y^{3}}{(1-Y)^{3}}\right)^{n}-\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{n}\left(\frac{XY^{3}}{(1-Y)^{3}}\right)^{n}-\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{n}\left(\frac{XY}{1-Y}\right)^{n}$
$=-\sum_{n=1}^{\infty}\frac{X^{n}Y^{2n}\bigl{(}3+(X-2)Y\bigr{)}^{n}}{n}\sum_{j=0}^{\infty}\binom{3n+j-1}{3n-1}Y^{j}$
$\quad+\sum_{n=1}^{\infty}\frac{(-1)^{n}X^{n}Y^{3n}}{n}\sum_{j=0}^{\infty}\binom{3n+j-1}{3n-1}Y^{j}+\sum_{n=1}^{\infty}\frac{(-1)^{n}X^{n}Y^{n}}{n}\sum_{j=0}^{\infty}\binom{n+j-1}{n-1}Y^{j}$
$=-\sum_{n=1}^{\infty}\frac{X^{n}Y^{2n}}{n}\sum_{k=0}^{n}\binom{n}{k}3^{n-k}(X-2)^{k}Y^{k}\sum_{j=0}^{\infty}\binom{3n+j-1}{3n-1}Y^{j}$
$\quad+\sum_{n=1}^{\infty}\frac{(-1)^{n}X^{n}}{n}\sum_{\nu=3n}^{\infty}\binom{\nu-1}{3n-1}Y^{\nu}+\sum_{n=1}^{\infty}\frac{(-1)^{n}X^{n}}{n}\sum_{\nu=n}^{\infty}\binom{\nu-1}{n-1}Y^{\nu}$
$=-\sum_{\nu=2}^{\infty}Y^{\nu}\sum_{\mu=1}^{\infty}\sum_{k=0}^{\mu}\frac{X^{\mu}}{\mu}\binom{\mu}{k}3^{\mu-k}\sum_{l=0}^{k}\binom{k}{l}X^{j}(-2)^{k-l}\binom{\nu+\mu-k-1}{3\mu-1}$
$\quad+\sum_{\nu=3}^{\infty}Y^{\nu}\sum_{m=1}^{\left\lfloor\frac{\nu}{3}\right\rfloor}\binom{\nu-1}{3m-1}\frac{(-1)^{m}X^{m}}{m}+\sum_{\nu=1}^{\infty}Y^{\nu}\sum_{m=1}^{\nu}\binom{\nu-1}{m-1}\frac{(-1)^{m}X^{m}}{m}$
$=-\sum_{\nu=2}^{\infty}Y^{\nu}\sum_{m=1}^{\infty}X^{m}\sum_{l=0}^{\left\lfloor\frac{m}{2}\right\rfloor}\sum_{k=l}^{m-l}\frac{3^{m-l-k}(-2)^{k-l}}{m-l}\binom{m-l}{k}\binom{k}{l}\binom{\nu+m-l-k-1}{3m-3l-1}$
$\quad+\sum_{n=3}^{\infty}Y^{n}\sum_{m=1}^{\left\lfloor\frac{n}{3}\right\rfloor}\binom{n-1}{3m-1}\frac{(-1)^{m}X^{m}}{m}+\sum_{n=1}^{\infty}Y^{n}\sum_{m=1}^{n}\binom{n-1}{m-1}\frac{(-1)^{m}X^{m}}{m}\,.$
Comparing the coefficients, we have
$\prod_{i=1}^{3}(\alpha_{i}^{n}-1)$
$=n\sum_{m=0}^{\left\lfloor\frac{n}{3}\right\rfloor-1}\binom{n-1}{3m+2}\frac{(-1)^{m+1}X^{m+1}}{m+1}+n\sum_{m=0}^{n-1}\binom{n-1}{m}\frac{(-1)^{m+1}X^{m+1}}{m+1}$
$\quad-n\sum_{m=0}^{\infty}X^{m+1}\sum_{l=0}^{\left\lfloor\frac{m+1}{2}\right\rfloor}\sum_{k=l}^{m-l+1}\frac{3^{m-l-k+1}(-2)^{k-l}}{m-l+1}\binom{m-l+1}{k}\binom{k}{j}\binom{n+m-l-k}{3m-3l+2}\,.$
Thus, by (16) we have
$n^{2}$ $=\left(\sum_{m=0}^{n-1}\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,3)X^{m}\right)\biggl{(}\sum_{m=0}^{\left\lfloor\frac{n}{3}\right\rfloor-1}\binom{n-1}{3m+2}\frac{(-1)^{m}X^{m}}{m+1}+\sum_{m=0}^{n-1}\binom{n-1}{m}\frac{(-1)^{m}X^{m}}{m+1}$
$\qquad+\sum_{m=0}^{\infty}X^{m}\sum_{l=0}^{\left\lfloor\frac{m+1}{2}\right\rfloor}\sum_{k=l}^{m-l+1}\frac{3^{m-l-k+1}(-2)^{k-l}}{m-l+1}\binom{m-l+1}{k}\binom{k}{l}\binom{n+m-l-k}{3m-3l+2}\biggr{)}\,.$
Hence, we have the following.
Theorem 4.
For $11\leq m\leq\left\lfloor\frac{n}{3}\right\rfloor-$ 1, we have
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,3)$
$=-\sum_{j=0}^{m-1}\frac{1}{n^{2}}\biggl{(}\sum_{l=0}^{\left\lfloor\frac{m-j+1}{2}\right\rfloor}\sum_{k=l}^{m-j-l+1}\frac{3^{m-j-l-k+1}(-2)^{k-l}}{m-j-l+1}\binom{m-j-l+1}{k}\binom{k}{l}\binom{n+m-j-l-k}{3m-3j-3l+2}$
$\qquad+\left(\binom{n-1}{3m-3j+2}+\binom{n-1}{m-j}\right)\frac{(-1)^{m-j}}{m-j+1}\biggr{)}\mathfrak{Z}_{n}^{\star}(\zeta_{n};j,3)$
with $\mathfrak{Z}_{n}^{\star}(\zeta_{n};0,3)=1$.
By applying this formula repeatedly, we can get explicit expressions of $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,3)$ ($m=1,2,3,4$) as follows.
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};1,3)$ $=-\frac{(n-1)(n-3)}{8}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};2,3)$ $=-\frac{(n+1)(n-1)(n^{4}-650n^{2}+3780n-5291)}{12\cdot 7!}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};3,3)$ $=\frac{(n+1)(n-1)(n-3)(8n^{5}+4n^{4}-2005n^{3}+6985n^{2}+11357n-36509)}{60\cdot 8!}\,,$
$\mathfrak{Z}_{n}^{\star}(\zeta_{n};4,3)$ $=\frac{(n+1)(n-1)}{2\cdot 15!}(703n^{10}-1109042n^{8}+4324320n^{7}+165698599n^{6}-1101079980n^{5}$
$\quad-545336011n^{4}+15460525080n^{3}-18165129202n^{2}-47055628620n+76385318153)\,.$
Acknowledgement
This work was partly supported by JSPS KAKENHI Grant Number 24K22835.